I got help with task 1 and 2 . can you help me with task 3 and 4 of this question. please help me step for step thanks.

A signal x[n] modulated by multiplying it by a carrier wave cos(2*p1/cm) to form the signal z[n] = cos(2p1Vcm)x[n] ·The mo
Answer #1

2 2- 42- =-o-6 12Ł = 0.6

we chooe HV) as 〉レ VeVe